OpenStudy (anonymous):

An exam consists of 15 true-or-false questions. If a student guesses at every answer, what is the probability that he or she will answer exactly 9 questions correctly? (Round your answer to three decimal places.)

OpenStudy (anonymous):

1/2(9/15)

OpenStudy (anonymous):

You can simply use the binomial distribution. In this case, you are calculating for P(X = 9). So, (15 C 9) (.5^9)(.5^6)

OpenStudy (anonymous):

if there is exactly 9 right then the probability of that is .5^9 (50% chance of getting it right) times .5^6 because if you get exactly 9 then you have to miss 6 questions. Then multiply by 15 C 9 (15 choose 9) which basically means how many different ways can 9 things be arranged out of 15 things

Directrix (directrix):

C(15,9) * (1/2)^9 * (1/2)^6 = 5005/32768 = .153 approx